OpenStudy (anonymous):

My teacher gave us the answers, but he wants us to show the work. And I'm stuck.

OpenStudy (lukecrayonz):

ask away

OpenStudy (anonymous):

-n-8+4n=28 is my problem and my answer is 12, but how do I show the work?

OpenStudy (lukecrayonz):

-n-8+4n=28 Start by combining like terms, the n's, and end with -8+3n=28. Then, you want to isolate n. Move 8 by adding since it is a negative. 3n=36. Now, divide 36 by 3, which is 12, and that is your answer, n=12.

OpenStudy (anonymous):

With problems like this your whole objective is to isolate your n variable on one side and the constants on the other side. so using the properties of multiplication, division, addition, and subtraction, you try to get n on one side and constantson the other

OpenStudy (lukecrayonz):

I'll draw the work for you to copy.

OpenStudy (lukecrayonz):